OpenStudy (anonymous):

In an AP , if the sum of 4th and 8th term is 70 and its 15th term is 80.Find the sum of first 25 terms

OpenStudy (king):

is common diff. given

OpenStudy (amistre64):

so what is it we know: spose we start at a4, or some other term and call it a1 then: a1 + a5 = 70 a12 = 80 but not sure if this thought is good yet

OpenStudy (amistre64):

not alot to go on :/

OpenStudy (amistre64):

it would help if we knew what kind of sequence it was, geometric, artithmetic, other

OpenStudy (king):

its given it is a arithmetic progression

OpenStudy (amistre64):

ahh, AP. I thought that meant it was an AP class lol

OpenStudy (amistre64):

1, 69 in 5 terms would give us a5 = a1 + d(5-1) 69 = 1 + 4d 68/4 = d = 17 1, 18, 35, 52, 69 is a way to find any d how far off the mark are we? a12 = 1 + 17(11) 80 = 1 + 181 ..... a bit ff, but it narrows it down

OpenStudy (king):

dude i think we first need to find comm. diff

OpenStudy (amistre64):

may 30 and 40? 40 = 30 + 4d 10/4 = d = 5/2 80 = 30 + (5/2)11 80 = 30 + 55/2 80 = 57 is close :)

OpenStudy (king):

don't use trial and error method.....

OpenStudy (amistre64):

finding a common diff is the way to go, but trial and error is a good way to narrow down the options

OpenStudy (king):

its not good

OpenStudy (amistre64):

im not saying its the best option; but its not an unusable option

OpenStudy (king):

ya but then we should do sumthin else

OpenStudy (king):

(Think)(Think)(Think)(Think)(Think)(Think)

OpenStudy (amistre64):

exactly; get a feel for it, then let your mind expand upon it ;)

OpenStudy (king):

see a4=a+3d,a5=a+4d so a4+a5=2a+7d 2a+7d=70 now what?

OpenStudy (amistre64):

you know at least one term, sub it in for value someplace

OpenStudy (king):

and since 15th term =80, a+14d=80

OpenStudy (amistre64):

a4 + a5 is a bit off; try a4 + a8 = 70 maybe?

OpenStudy (king):

a=20,d=30/7

OpenStudy (king):

we can now find all nos and add them

OpenStudy (amistre64):

a4 = a1 + 3d a8 = a1 + 7d ------------ 70 = 2a1 + 10d

OpenStudy (anonymous):

a+3d+a+7d=70 2a+10d=70 2(a+5d)=70 a+5d=35-------1 a+14d=80------2 subtract 1 from 2 a+14d=80 -a-5d=-35 ----------- 9d=45 d=5 substituting d in 1 a+5(5)=35 a+25=35 a=10 Sn=(n/2)*[2a+(n-1)d] S25=(25/2)*[(2*10)+(24)5] =(25/2)*[20+120] =(25/2)*(140) =1750

OpenStudy (amistre64):

from 1 and 2 we can cramer it to get a and d as well: d = 80-35 / 14-5 = 45/9 = 5 a = 80(5)-14(35) /-9 = -90 /-9 = 10